# From Many Models, One: Macroeconomic Forecasting with Reservoir Ensembles

### v1

@@ Abstract
Model combination is a powerful approach to achieve superior performance with a set of models than by just selecting any single one. We study both theoretically and empirically the effectiveness of ensembles of Multi-Frequency Echo State Networks (MFESNs), which have been shown to achieve state-of-the-art macroeconomic time series forecasting results (Ballarin et al., 2024a, ). Hedge and Follow-the-Leader schemes are discussed, and their online learning guarantees are extended to the case of dependent data. In applications, our proposed Ensemble Echo State Networks show significantly improved predictive performance compared to individual MFESN models.
@@ 1 Introduction
Many mathematical, statistical, and econometric theories have been developed to solve the problem of predicting stochastic (and deterministic) processes. Yet, in empirical settings, it is often hard to judge which model – or, even, which model class – among a large pool of potential candidates most adequately fits the task at hand. A model that performs best most of the time may fail during some periods when a different prediction scheme becomes more adequate. This phenomenon is pronounced in forecasting problems, for example, involving macroeconomic time series. Structural breaks, financial crises, and policy regime shifts can significantly affect the performance of forecasting models. Since the relative performance of individual models can vary substantially over time, it is natural, when model updating is infeasible or unsuitable, to revise the choice of model(s) as new information becomes available.
@@
The model combination approach formalizes the idea of incorporating different prediction schemes. In this framework, the forecaster relies on a finite set, called the ensemble, of distinct models/algorithms, termed experts, to predict an outcome of interest. The forecaster scores the experts according to their predictive accuracy over time based on their losses and iteratively reweights them. Thus, ensemble methods can adjust to changes in time series dynamics, favoring distinct models across different periods. The objective is to construct a sequence of data-driven weights that optimally aggregate individual expert forecasts to minimize cumulative loss over time. A well-designed combination scheme is expected to robustly select the best model(s) within the ensemble and closely match the minimum loss if it were known in hindsight. The ensemble framework can also be considered from a game-theoretic perspective: The forecaster competes against an adversary who generates outcomes to maximize regret, hence motivating schemes that perform well even in worst-case scenarios.
In this paper, we bridge the rich literature on online ensemble learning (see for example the overview by Shalev-Shwartz,, 2012) with that of econometric forecast combinations (Timmermann,, 2006) in a general time series setup. From a theoretical point of view, we study the properties of Follow-the-Leader (FTL) and Decreasing Hedge algorithms in the stochastic setting under the assumptions of i.i.d. (across time) or $\varphi$-mixing losses by building on the developments in Mourtada and Gaïffas, (2019) and de Rooij et al., (2014), respectively. We derive new Hoeffding- and Bernstein-type regret bounds for both algorithms, showing that the maximum error a forecaster can make – in terms of cumulative expected regret, which we formally define in Section 3 – is upper bounded by an explicit constant. While the dependence on the number of experts is mild (logarithmic), the gap between the least expected loss and the runner-up plays a key role in our bounds. Then, we build upon previous work Ballarin et al., 2024a to demonstrate the performance gains offered by ensembles that combine Multi-Frequency Echo State Network (MFESN) models, which were shown to have higher accuracy than state-of-the-art methods such as MIxed DAta Sampling (MIDAS) (Ghysels et al.,, 2004, 2007) and Dynamic Factor Models (DFM) (Geweke,, 1977; Sargent and Sims,, 1977). The MFESN model class introduced in Ballarin et al., 2024a is based on Echo State Networks (ESNs), which, at a high level, are recurrent-type neural networks with state parameters that are fixed at random initializations (Jaeger,, 2010; Maass,, 2011; Lukoševičius and Jaeger,, 2009). This model family has been actively explored in terms of their approximation, generalization, and memory properties (Grigoryeva and Ortega, 2018b, ; Grigoryeva and Ortega, 2018a, ; Gonon and Ortega,, 2020; Gonon et al., 2020b, ; Gonon et al.,, 2023; Ballarin et al., 2024b, ). Our empirical study combines the MFESN framework with ensembles for the prediction of quarterly U.S. GDP growth utilizing the medium-sized set of monthly and daily financial and macroeconomic variables originally compiled in Ballarin et al., 2024a . We find that, by constructing ensembles of MFESNs instantiated with different random parameter draws, we can substantially reduce mean-squared forecasting errors. When we additionally change a model hyperparameter (the state leak rate) across the ensemble, improvements upwards of $40\%$ against the baseline MFESNs in Ballarin et al., 2024a can be achieved. These results set new benchmarks in the prediction of U.S. GDP growth with multi-frequency regressors. Finally, we assess the evolution of the ensemble weights in our forecasting exercise, showing that a handful of models receive the most weight.
Forecast combination is a long-standing idea in econometrics and time series forecasting, well in development already half a century ago (Bates and Granger,, 1969; Winkler and Makridakis,, 1983; Clemen,, 1989; Granger,, 1989). [^2] Boosting, bagging, and random forest methods, which explicitly combine many learners into one predictor to improve performance, have also become increasingly popular (Athey,, 2019; Athey and Imbens,, 2019). Some more recent examples include, for example, Hansen, (2008); Cheng and Hansen, (2015); Kim and Swanson, (2018); den Reijer and Johansson, (2019); Bolhuis and Rayner, (2020); Fuleky, (2020); Yoon et al., (2011); Masini et al., (2023). We take a frequentist viewpoint, as Bayesian combination schemes are often computationally expensive (Wang et al.,, 2023). Linear combination methods, where the forecaster’s prediction is a weighted sum of the experts’ outputs, have been classically considered, but nonlinear and shrinkage approaches are also possible (Timmermann,, 2006). In this work, we concentrate on the former, as it is the most widely used setup in the literature on online ensemble learning. An important strand of interest concerns the combination of density forecasts (Wallis,, 2011). However, since this work focuses on mean-squared error prediction, we leave this aspect, as well as the question of uncertainty quantification, open for future work. We emphasize that the approach we take in this paper is quite different from that often considered in econometrics. First, we do not view the modeling combination problem as a question of estimating weights (see, for example, Timmermann,, 2006, Section 3), the complexity of which is often cited as the reason simple schemes, like averaging, work best. In fact, our theoretical analysis does not require any analysis of the correlation structure between forecasts (and their losses). The online weighting schemes we consider are time-varying, meaning they can adapt over time at the cost of not potentially incurring an early sub-optimality cost. Moreover, the empirical success of a single model scheme like Follow-the-Leader is not in general comparable to simply choosing the single best model. In fact, FTL selects the forecast of the model committing the smallest observed cumulative loss, which may change every period. Our proposed application of online ensemble schemes is thus aligned with the objective of “keeping combinations sophisticatedly simple(Wang et al.,, 2023). Finally, our empirics focus on using ensembles consisting of distinct realizations of a single model class: We view this as an effective and efficient way to guard against the luck-of-the-draw inherent in random-parameter models.
@@
The literature on online combination schemes and ensembles is rich. In the standard framework of prediction with expert advice (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997; Vovk,, 1995; Cesa-Bianchi and Lugosi,, 2006), or its variant – the Hedge setting – much of recent work has been concentrated on developing combination algorithms that enjoy (desirably sharp) guarantees on regret, defined as the difference between the cumulative loss of the forecaster and that of the best among $K\geq 1$ experts. Follow-the-Leader directly assigns uniform weights at each round to the (subset of) model(s) with the smallest cumulative loss so far, and is known to guarantee an $O(\log T)$ regret for strongly convex losses in the stochastic setting (with $T$ a time horizon). While FTL is known to perform extremely well with stochastic losses (van Erven et al.,, 2011), it also demonstrates poor performance for worst-case data. The Hedge algorithm (Freund and Schapire,, 1997), also called the exponentially weighted averaged forecaster, offers a smoothed version of the FTL scheme: Under optimal learning rate, it enjoys an anytime worst-case $O(\sqrt{T\log K})$ regret bound. However, in the stochastic setting, one should allow the learning rate of Hedge to be data-adaptive to achieve regret comparable with FTL. This idea is behind the so-called Adaptive Hedge (AdaHedge) scheme (van Erven et al.,, 2011; de Rooij et al.,, 2014). Significant effort has been made to develop strategies combining the best of the two approaches, see, for example, Koolen et al., (2014); Wintenberger, (2017); Mourtada and Gaïffas, (2019); Ito et al., (2024); Wintenberger, (2024).
@@ 2 Preliminaries
We begin by introducing the general econometric setting of interest, which involves constructing predictions of a time series by combining (or selecting) models from a given set of experts.
@@ 2.1 General Setup
Let $\{Y_{t}\}_{t\in\mathbb{Z}}$, $Y_{t}\in\mathbb{R}$, be the target time series, and $\{Z_{t}\}_{t\in\mathbb{Z}}$, $Z_{t}\in\mathbb{R}^{d}$, a vector of regressors, which may include lagged values of $Y_{t}$. We consider the problem of predicting the target series $Y_{t+h}$ at the forecasting horizon $h\in\mathbb{N}^{+}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a fixed probability space on which all the random variables are defined. In the most general setting the law $\mathbb{P}_{t}$ of $(Y_{t},Z_{t})$ may vary with $t$. For each $t\in[T]$, let $Z_{t}$ be $\mathcal{F}_{t}$-measurable with $\mathcal{F}_{t}:=\sigma\left(Z_{t}\right)$. For $h\in[H]$, the $h$-step-ahead minimum mean square error (MSE) predictor of $Y_{t+h}$ at time $t$ is the conditional expectation function (CEF) given by
@@
Two challenges arise even in the one-step-ahead prediction setting ($h=1$) when solving (1): The law $\mathbb{P}_{t}$ of $(Y_{t},Z_{t})$ may be time-varying (nonstationary), and minimizing MSE over (possibly infinite-dimensional) $\mathcal{G}$ class of functions can be infeasible or computationally and statistically demanding. The latter problem is typically addressed by structural assumptions on the joint data-generating process $(Y_{t},Z_{t})$. For example, standard choices include linear models (Brockwell and Davis,, 2006), semi- or nonparametric classes of sufficient regularity (Stone,, 1982; Li and Racine,, 2009; Tsybakov,, 2009), or a reproducing kernel Hilbert space setting (Christmann and Steinwart,, 2008; Schölkopf et al.,, 2002; Berlinet and Thomas-Agnan,, 2011). The first issue is more subtle, as classical prediction settings often assume independent or stationary time series data. [^3] One may also want to work under local stationarity (Dahlhaus,, 1997, 2012) or non-stationarity assumptions, although arguably these setups would require significantly more complex proof techniques. When $\mathbb{P}_{t}$ can evolve with $t$ without strong restrictions, the predictor in (1) must be updated over time, leading to an online learning formulation. In this work, we study a standard online aggregation setting in which multiple models (“experts”) are combined, each potentially excelling under different realizations of $\mathbb{P}_{t}$.
@@ 2.2 Prediction with Experts
The concept of prediction with experts goes back to Littlestone and Warmuth, (1994); Foster, (1991); Foster and Vohra, (1993); Vovk, (1995), and Cesa-Bianchi et al., (1997). A forecaster is interested in obtaining at each time step $t\in[T]$ a prediction $\widehat{Y}_{t+h}$, for a horizon $h\in[H]$, based on $K\geq 1$ experts indexed by $k\in[K]$, each providing their own forecast $\widehat{Y}_{t+h}^{(k)}$. The forecaster constructs a convex combination of $K$ expert predictions using a weight vector $\bm{\omega}_{t,h}:=(\omega^{(1)}_{t,h},\ldots,\omega^{(K)}_{t,h})^{\top}\in\mathbb{R}^{K}$ such that
@@
We will also use $\Delta^{K-1}:=\left\{\bm{\omega}\in\mathbb{R}_{+}^{K}:\bm{\omega}^{\top}\bm{1}_{K}=1\right\}$.
The accuracy of each forecast is assessed using a loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$. We define for each expert $k\in[K]$ and each $t\in[T]$ the associated loss at $h$-horizon forecasting task of predicting ${Y}_{t+h}$, $h\in[H]$, as
@@
the vector containing the losses of all $K$ experts as $\bm{\ell}_{t,h}:=(\ell^{(1)}_{t,h},\ldots,\ell^{(K)}_{t,h})^{\top}$ and the forecaster’s loss $\overline{\ell}_{t,h}:=\bm{\omega}_{t,h}^{\top}\bm{\ell}_{t,h}$. We also define the cumulative loss of expert $k$ and of the forecaster, incurred by time $t$ when forecasting $h$ steps ahead as
@@
The definition of $\varphi$-mixing dependence in (ii) and (iii) is in the sense of Definition 2.1 in Rio, (2017). The dependence conditions are imposed over time for experts, but no independence is assumed across $K$ experts.
@@
Assumption 1 is standard in online learning problems and algorithms and, up to a fixed rescaling, amounts to requiring the losses to be bounded. Assumption 2 imposes conditions on the realization of loss vectors over time, as the prediction algorithm evolves. Much of the literature focuses on Assumption 2(i), since i.i.d. losses are more straightforward to study within the statistical learning toolbox. For our purposes, however, mixing conditions such as Assumption 2(ii) are empirically more relevant, as they accommodate time series settings. Notice also that for any $k\in[K]$, since $0\leq\varphi_{n}^{(k)}\leq 1$, Assumption 4(i) implies Assumption 3(i) and Assumption 4(ii) implies Assumption 3(ii), respectively. Moreover, in each of these assumptions, part (ii) implies part (i).
Importantly, these assumptions are imposed on the loss processes $\{\ell^{(k)}_{t,h}\}_{t}$ (for each expert $k$ and horizon $h$), not necessarily on the joint data process $(Y_{t},Z_{t})$. In particular, if experts are fixed and the underlying process is mixing, then losses $\{\ell^{(k)}_{t,h}\}_{t}$ inherit mixing structure under measurable maps. Even under mild nonstationarity (for example, local stationarity), these incurred losses often remain weakly dependent, making uniform mixing a plausible condition.
In the following sections, we restrict our analysis to the case of a single forecasting horizon of interest $h\in\mathbb{N}$ and hence use or drop the $h$ subscript where convenient (for example, we write $\ell_{t}^{(k)}$ to denote $\ell_{t,h}^{(k)}$). Our results apply naturally to the multi-horizon case by analyzing cumulative regret separately for each horizon $h\in[H]$.

### v2

@@ Abstract
Model combination is a powerful approach for achieving superior performance compared to selecting a single model. We study both theoretically and empirically the effectiveness of ensembles of Multi-Frequency Echo State Networks (MFESNs), which have been shown to achieve state-of-the-art macroeconomic time series forecasting results (Ballarin et al., 2024a, ). The Hedge and Follow-the-Leader schemes are discussed, and their online learning guarantees are extended to settings with dependent data. In empirical applications, the proposed Ensemble Echo State Networks demonstrate significantly improved predictive performance relative to individual MFESN models.
@@ 1 Introduction
Many mathematical, statistical, and econometric tools have been developed to solve the problem of predicting stochastic and deterministic processes. In empirical applications, however, it is often difficult to identify which model, or even which model class, is most appropriate for a given task. Models that perform well on average may fail during particular periods. This situation is common in forecasting settings, for example, involving macroeconomic time series. Structural breaks, financial crises, and policy regime shifts can significantly affect predictive accuracy. Since forecasting performance can vary substantially over time, it is natural, when model updating is infeasible or unsuitable, to revise the choice of model(s) as new information becomes available.
@@
The model combination approach formalizes the use of multiple prediction schemes. The forecaster relies on a finite set, referred to as an ensemble, of distinct models or algorithms, termed experts, to predict an outcome of interest. Experts in an ensemble are evaluated and reweighted over time based on their predictive losses. Ensemble methods can hence adapt to changing time series dynamics by favoring different models across periods. The objective is to construct data-driven weights that aggregate individual expert forecasts to minimize cumulative loss over time. A well-designed combination scheme is expected to select the best-performing models, achieving performance close to the minimum loss in hindsight.
Forecast combination is a long-standing idea in econometrics and time series analysis (Bates and Granger,, 1969; Winkler and Makridakis,, 1983; Clemen,, 1989; Granger,, 1989) [^2] Boosting, bagging, and random forest methods, which explicitly combine many learners into one predictor to improve performance, have also become increasingly popular (Athey,, 2019; Athey and Imbens,, 2019)., with a large literature documenting its empirical effectiveness in a variety of macroeconomic and financial forecasting contexts. Subsequent contributions have studied both linear and nonlinear combination schemes, as well as the role of time variation in forecast weights (see Timmermann,, 2006, for a comprehensive survey). More recent work has emphasized adaptive and data-driven approaches to forecast combination in high-dimensional and potentially unstable environments, including settings with structural change and model misspecification (e.g., Hansen, (2008); Cheng and Hansen, (2015); Kim and Swanson, (2018); den Reijer and Johansson, (2019); Bolhuis and Rayner, (2020); Fuleky, (2020); Masini et al., (2023)), as well as considered Bayesian (Wang et al.,, 2023) and density combination schemes (Wallis,, 2011).
@@
The literature on online combination schemes and ensembles is rich. In the standard framework of prediction with expert advice (Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997; Vovk,, 1995; Cesa-Bianchi and Lugosi,, 2006), or its variant known as the Hedge setting, much of the work focuses on developing combination algorithms that enjoy guarantees on regret, defined as the difference between the cumulative loss of the forecaster and that of the best among $K$ experts. Prominent examples include Follow-the-Leader (FTL), which assigns uniform weights at each round to the model(s) with the smallest cumulative loss so far; it is known to achieve an $O(\log T)$ regret for strongly convex losses in stochastic settings (with $T$ a time horizon), while potentially performing poorly under worst-case data (van Erven et al.,, 2011). The Hedge algorithm (Freund and Schapire,, 1997), also called the exponentially weighted forecaster, provides a smoothed alternative to FTL and enjoys an anytime worst-case $O(\sqrt{T\log K})$ regret bound under an optimal learning rate. In stochastic environments, allowing the learning rate of Hedge to be data-adaptive yields regret comparable to FTL, motivating schemes such as Adaptive Hedge (van Erven et al.,, 2011; de Rooij et al.,, 2014). Significant effort has been devoted to develop strategies that combine the best properties of these approaches (see, for example, Koolen et al., (2014); Wintenberger, (2017); Mourtada and Gaïffas, (2019); Ito et al., (2024); Wintenberger, (2024)).
In this paper, we bridge the literature on online ensemble learning (see, e.g., Shalev-Shwartz,, 2012) with econometric forecast combination (Timmermann,, 2006) in a general time series setup. From a theoretical perspective, we study the properties of Follow-the-Leader and Decreasing Hedge (DecHedge) algorithms in the stochastic setting for i.i.d. (in time) and $\varphi$-mixing losses, building upon the developments in Mourtada and Gaïffas, (2019) and de Rooij et al., (2014), respectively. For both algorithms, we derive Hoeffding- and Bernstein-type regret bounds that characterize their cumulative performance relative to the best expert in hindsight. While the dependence on the number of experts is mild (logarithmic), the gap between the least expected loss and the runner-up plays a key role in our bounds.
On the empirical side, we construct our ensembles using Multi-Frequency Echo State Network (MFESN) models introduced in Ballarin et al., 2024a , which have been shown to outperform state-of-the-art mixed-frequency forecasting methods such as MIxed DAta Sampling (MIDAS) (Ghysels et al.,, 2007) and Dynamic Factor Models (DFMs) (Geweke,, 1977; Sargent and Sims,, 1977). The MFESN models are based on Echo State Networks (ESNs), a class of recurrent-type neural networks with randomly sampled state parameters (Jaeger,, 2010; Maass,, 2011; Lukoševičius and Jaeger,, 2009). This model family has been actively studied in terms of approximation, generalization, and memory properties (Grigoryeva and Ortega, 2018b, ; Grigoryeva and Ortega, 2018a, ; Gonon and Ortega,, 2020; Gonon et al., 2020b, ; Gonon et al.,, 2023; Ballarin et al., 2024b, ). Using quarterly U.S. GDP growth and a set of $33$ monthly and daily financial and macroeconomic variables in Ballarin et al., 2024a , we show that MFESN ensembles substantially reduce mean-squared forecasting errors. Additionally, allowing for heterogeneity in random parameter draws and state leak rates across the ensemble yields up to $40\%$ improvements relative to baseline MFESNs. These results set new benchmarks in the prediction of U.S. GDP growth with multi-frequency regressors. Finally, we report the experts’ weight evolution in our forecasting exercise, showing that a handful of models receive most of the weight.
@@ 2 Preliminaries
In this section, we introduce the general econometric setting of interest, which involves constructing time series predictions by combining (or selecting) models from a set of experts.
@@ 2.1 General Setup
Let $\{Y_{t}\}_{t\in\mathbb{Z}}$, $Y_{t}\in\mathbb{R}$, be the target time series, and $\{Z_{t}\}_{t\in\mathbb{Z}}$, $Z_{t}\in\mathbb{R}^{d}$, a vector of regressors, which may include lagged values of $Y_{t}$. We consider the problem of predicting the target series $Y_{t+h}$ at the forecasting horizon $h\in\mathbb{N}^{+}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a fixed probability space on which all the random variables are defined. For each $t\in[T]$, let $Z_{t}$ be $\mathcal{F}_{t}$-measurable with $\mathcal{F}_{t}:=\sigma\left(Z_{t}\right)$. Under relatively weak conditions of the joint law of $Y_{t}$ and $Z_{t}$ the $h$-step-ahead minimum mean square error (MSE) predictor of $Y_{t+h}$ at time $t$, with $h\in[H]$, is the conditional expectation function (CEF) given by
@@
The core challenge arising when solving (1), even in the one-step-ahead prediction setting ($h=1$), is that minimizing MSE over (possibly infinite-dimensional) $\mathcal{G}$ class of functions can be infeasible or computationally and statistically demanding. This is typically addressed by structural assumptions on the joint data-generating process $(Y_{t},Z_{t})$. For example, standard choices include linear models (Brockwell and Davis,, 2006), semi- or nonparametric classes of sufficient regularity (Stone,, 1982; Li and Racine,, 2009; Tsybakov,, 2009), or a reproducing kernel Hilbert space setting (Christmann and Steinwart,, 2008; Schölkopf et al.,, 2002; Berlinet and Thomas-Agnan,, 2011). In this work, we study a standard online aggregation setting in which multiple models (“experts”) are combined, each potentially excelling under different finite sample realizations, but always assuming that a best-performing model exists (see Section 3). The focus of the paper is on studying schemes to weight experts in order to produce a combined forecast that over time aligns with the optimal expert.
@@ 2.2 Prediction with Experts
The concept of prediction with experts goes back to the works of Littlestone and Warmuth, (1994); Foster, (1991); Foster and Vohra, (1993); Vovk, (1995); Cesa-Bianchi et al., (1997). A forecaster is interested in obtaining, at each time step $t\in[T]$, a prediction $\widehat{Y}_{t+h}$, for a horizon $h\in[H]$, based on $K\geq 1$ experts indexed by $k\in[K]$, each providing their own forecast $\widehat{Y}_{t+h}^{(k)}$. The forecaster constructs a convex combination of $K$ expert predictions using a weight vector $\bm{\omega}_{t,h}:=(\omega^{(1)}_{t,h},\ldots,\omega^{(K)}_{t,h})^{\top}\in\mathbb{R}^{K}$ such that
@@
We will also use $\Delta^{K-1}:=\{\bm{\omega}\in\mathbb{R}_{+}^{K}:\bm{\omega}^{\top}\bm{1}_{K}=1\}$ to indicate the real simplex.
The accuracy of each forecast is assessed using a loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$. We define, for each expert $k\in[K]$ and each $t\in[T]$, the associated loss at $h$-horizon forecasting task of predicting ${Y}_{t+h}$, $h\in[H]$, as
@@
the vector containing the losses of all $K$ experts as $\bm{\ell}_{t,h}:=(\ell^{(1)}_{t,h},\ldots,\ell^{(K)}_{t,h})^{\top}$ and the forecaster’s loss $\overline{\ell}_{t,h}:=\bm{\omega}_{t,h}^{\top}\bm{\ell}_{t,h}$. We also define the cumulative loss of expert $k$ and of the forecaster, incurred by time $t$ when forecasting $h$ steps ahead, as
@@
The definition of $\varphi$-mixing dependence in (ii) and (iii) is in the sense of Definition A.1 in the Appendix, following Rio, (2017). The dependence conditions are imposed on losses (hence, data and experts) over time, but no independence is assumed across the $K$ losses.
@@
Assumption 1 is standard in online learning and, up to a fixed rescaling, requires losses to be bounded. Assumption 2 imposes conditions on the temporal dependence of loss vectors, as the prediction algorithm evolves. While much of the literature focuses on the i.i.d. case, mixing conditions such as Assumption 2(ii) are empirically more relevant in time-series applications. Notice also that for any $k\in[K]$, since $0\leq\varphi_{n}^{(k)}\leq 1$, Assumption 4(i) implies Assumption 3(i) and Assumption 4(ii) implies Assumption 3(ii). Moreover, in each of these assumptions, part (ii) implies part (i). All assumptions are imposed on the loss processes $\{\ell^{(k)}_{t,h}\}_{t}$, rather than on the joint data process $(Y_{t},Z_{t})$. If experts are fixed and $(Y_{t},Z_{t})$ is mixing, the induced loss processes inherit this property under measurable maps. Even under mild nonstationarity (e.g., local stationarity), these losses often remain weakly dependent, making uniform mixing a plausible assumption in practice.
In the remainder of the paper, we focus on a single forecasting horizon $h\in\mathbb{N}$ and hence drop the $h$ subscript where convenient. Our results extend directly to the multi-horizon case by analyzing cumulative regret separately for each horizon $h\in[H]$.